q1=+10 nC and q2=-5 nC are separated by 0.10 m distance. If a third charge q3=2 nC is placed along with the first two charges, where should it be so that the net electric force of q1 and q2 on q3 is zero (0)?


1
Expert's answer
2021-05-30T13:30:59-0400
"k\frac{q_1}{(x+d)^2}=k\frac{q_2}{x^2}\\\frac{10}{(x+0.1)^2}=\frac{5}{x^2}
\\x=0.24\ m"

It is the distance from the second charge.